40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3798     Entry time: Wed Oct 27 16:15:35 2010
Author: Suresh 
Type: Update 
Category: SUS 
Subject: Re-glued magnet to the PRM 

Thanh and I re-glued the magnet to the PRM following the procedure outlined by Jenne

The PRM in the gluing fixture has been placed in the little foil house and left to cure for a day.

If all goes well the balancing the PRM will be done tomorrow.

 

 

ELOG V3.1.3-